$SPAD/src/input leg.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool leg.output
)set message test on
)set message auto off
)clear all

———-

2 Legendre Polynomials

— * —

--S 1 of 4
p(n | n=0) == 1
--R
--R                                                                   Type: Void
--E 1
--S 2 of 4
p(n | n=1) == x
--R
--R                                                                   Type: Void
--E 2
--S 3 of 4
p(n | n>1) == ((2*n-1)*x*p(n-1)-(n-1)*p(n-2))/n
--R
--R                                                                   Type: Void
--E 3
--S 4 of 4
p 3
--R
--R   Compiling function p with type Integer -> Polynomial(Fraction(
--R      Integer))
--R   Compiling function p as a recurrence relation.
--R

--R        5  3   3
--R   (4)  - x  - - x
--R        2      2
--R                                          Type: Polynomial(Fraction(Integer))
--E 4
)spool
)lisp (bye)

———-

References^{[1] nothing}

5